OpenStudy (tay1lynn):

Evaluate [(40 + 5) − 32] ÷ 9 ⋅ 2. @Photon

OpenStudy (photon336):

alright so first you need to do what's in the ( )

OpenStudy (photon336):

@tay1lynn

OpenStudy (photon336):

so the first operation would be addition

OpenStudy (tay1lynn):

ok

OpenStudy (tay1lynn):

so it would be 45-32 /9-2?

OpenStudy (photon336):

\[\frac{ (45-32) }{ 9 }•2\]

OpenStudy (tay1lynn):

oh ok

OpenStudy (photon336):

\[\frac{ 13 }{ 9 }•2 \]

OpenStudy (tay1lynn):

I got 2.8

OpenStudy (photon336):

\[\frac{ 13 }{ 9 }•\frac{ 2 }{ 1 } = \frac{ 26 }{ 9 } = 2.8\]

OpenStudy (tay1lynn):

yay I got it right